OpenStudy (anonymous):

Find the constant a such that the function is continuous on the entire real line.

OpenStudy (anonymous):

\[f(x) = 3x ^{2} , x \ge1\] \[f(x) = ax - 4, x \le1\]

OpenStudy (zarkon):

7

OpenStudy (anonymous):

Thanks. How'd you find it?

OpenStudy (zarkon):

\[3(1)^2=3\] \[a(1)-4=3\Rightarrow a=7\]

OpenStudy (anonymous):

Ah, ok. Thanks